OpenStudy (gabylovesyou):

What are the solutions of 4x2 − 4x − 3 = 0? x = −1 over 2, x = 3 over 2 x = 1 over 2, x = −3 over 2 x = 1 over 4, x = −3 x = −1 over 4, x = 3

OpenStudy (anonymous):

here, Over means what?

OpenStudy (gabylovesyou):

\[x = -\frac{ 1 }{ 2 }, x = \frac{ 3 }{ 2 }\] \[x = \frac{ 1 }{ 2 }, x = - \frac{ 3 }{ 2 }\] \[x = \frac{ 1 }{ 4 }, x = -3\] \[x = - \frac{ 1 }{ 4 }, x = 3\]

OpenStudy (anonymous):

x=−1 /2 ,x=3/ 2

OpenStudy (precal):

use the quadratic formula

OpenStudy (anonymous):

4x^2-6x+2x-3=0 2x(2x-3)+1(2x-3)=0 (2x+1)(2x-3)=0 solve and get ans